the computational part isn't right. I've asked 2 times already and every answer has been wrong :(

A critical reaction in the production of energy to do work or drive chemical reactions in biological systems is the hydrolysis of adenosine triphosphate, ÄTP, to adenosine diphosphate, ADP, as described by ATPaq) +H which ΔGnn--30.5 kJ/mol at 37.0 ℃ and pH 7.0. Calculate the value of Δ which [ATPI-5.0 mM, [ADPI 0. in a biological cell in 20 mM, and IHPo21-5.0 mM. Number k.J/ mol Is the hydrolysis of ATP spontaneous under these conditions? O yes no

same for this one.
Answer #1

delG= delG0+ RTlnK

K= Equilibrium constant which is written as

K= [ADP] [HPO4-2]/ [ATP] = 5*5/0.2 =125*10-3M

delG= -30.5+8.314*(37+273.15)ln (125*10-3)/1000 Kj

=-30.5-5.36 =-35.86 Kj/mol

The delG suggests the process to be spontaneous.

b) for the second reaction N2(g) +3H2(g)----> 2NH3(g)

delG= delG0+ RTlnK, K is equilibrium constant

K = [NH3]2/ [N2] [H2]3 = (0.85)2/{0.15*(0.1)3}=5667

delG= -72.6+8.314*298.15*ln(5667)/1000= -72.6+2.48 =-70.12 Kj/mol